# The JWST-NIRCam View of Sagittarius C. III. The Extinction Curve

### v1

@@ 5.3.2 Other extinction curves
To compare our extinction curve with others derived from regions outside the GC, we selected the extinction curves reported by Indebetouw et al. (2005), Román-Zúñiga et al. (2007), Damineli et al. (2016), and Fahrion & De Marchi (2023), as they represent a variety of environments. Indebetouw et al. (2005) derived their extinction curve using 2MASS photometry and Spitzer data in the 3–8 m range, targeting two lines of sight: (a quiescent region) and (the giant H II region RCW 49). Román-Zúñiga et al. (2007) obtained their extinction curve for the 1.25–7.76 m range using ground-based observations with ESO NTT and VLT, as well as Spitzer-IRAC data, focusing on Barnard 59, a dense, star-forming core in the Pipe Nebula. Damineli et al. (2016) combined photometric and spectroscopic data from multiple instruments to derive the extinction curve for Westerlund 1 over the 0.4–4.8 m range. Finally, Fahrion & De Marchi (2023) derived the extinction curve for 30 Doradus in the Large Magellanic Cloud using JWST-NIRCam observations in the 0.3–4.8 m range.
@@
Figure 7 shows the comparison between our extinction curve and those of Indebetouw et al. (2005) and Román-Zúñiga et al. (2007) (left panel), as well as those of Damineli et al. (2016) and Fahrion & De Marchi (2023) (centre panel). In both cases, we included the J band extinction ratio by interpolating our extinction curve. For the left panel, we adopted the isophotal wavelengths of 2MASS and IRAC filters convolved with a K2 III stellar spectrum, following the approach in Indebetouw et al. (2005) and Román-Zúñiga et al. (2007). For the right panel, we used the J band effective wavelength provided by Damineli et al. (2016).
In addition, in the right panel of Fig. 7 we include theoretical extinction curves from Draine (2003) for and , and from Hensley & Draine (2023), as implemented in the dust_extinction package by Gordon (2024). These grain models are provided for reference only, and a detailed comparison is beyond the scope of this work.
@@
Table 5 presents the interpolated extinction ratios for the literature extinction curves, alongside those derived in this work, allowing for a more direct comparison. As seen in both the left panel of Fig. 5 and the table, our results are consistent with those reported by Indebetouw et al. (2005). This supports the hypothesis that the infrared extinction curve is independent of the line of sight, as their curve was derived from two distinct Galactic directions. Our extinction ratios also match those reported by Román-Zúñiga et al. (2007), although only in the SW regime. Both studies used Spitzer-IRAC data in the LW range; the discrepancies with our results may be explained by the broad Spitzer filter profiles and the fact that Barnard 59 is a cold cloud, where the presence of ices can significantly affect extinction in this regime (Riaz et al., 2009; Chu et al., 2020; Chu & Hodapp, 2021). Additionally, we computed the ratio using our extinction curve, as described in Sect. 5.3.1, obtaining a value of . This is higher than the values reported by Indebetouw et al. (2005) () and Román-Zúñiga et al. (2007) ().
@@
Our extinction curve also shows agreement with that reported by Damineli et al. (2016) in the SW regime—except for the ratio—but deviates in the LW regime, as illustrated in the right panel of Fig. 7 and in Table 5. However, a robust comparison is limited by the fact that Damineli et al. (2016) do not provide uncertainties for their extinction ratios. Unlike the B59 case, this discrepancy is unlikely to be driven by ice absorption, as the line of sight towards Westerlund 1 shows little evidence of ice (Guarcello et al., 2025). Similarly, our extinction curve agrees with that of Fahrion & De Marchi (2023), but the associated uncertainties are too large to allow for firm conclusions.
@@
While detailed modelling of the dust properties in Sgr C is beyond the scope of this work, we note that the theoretical curve corresponding to from Draine (2003) appears to be a closer match to our data than the model. This is consistent with expectations for denser environments with larger grains. Our results differ from the Hensley & Draine (2023) model in the LW regime. However, we caution that our extinction curve is derived from a localized region, and no strong conclusions about dust grain properties or can be drawn.
@@
In summary, our extinction curve follows the same general trend as the selected literature curves derived outside the GC. It shows excellent agreement with the curve by Indebetouw et al. (2005), reinforcing the idea of a broadly line-of-sight independent infrared extinction curve. It is also consistent with the curve by Fahrion & De Marchi (2023), although the large uncertainties prevent a definitive assessment.
@@ 6 Conclusions
In this work, we aimed to determine the extinction curve towards the Sgr C region using JWST-NIRCam data, and to evaluate its consistency with previous results. We applied two different methods based on the slopes of the RC features to derive the extinction ratios. The results are shown in Tables 2 and 3.
We found that the extinction curve flattens in the LW regime (3.6–4.8 m), in agreement with previous literature. Although we obtained different extinction indices for the SW and LW regimes, we found no significant wavelength dependence within each regime. However, using GALACTICNUCLEUS data overlapping with our JWST field, we derived an H– extinction index consistent with Nogueras-Lara et al. (2020), but not with our own results. We attribute this discrepancy mainly to the difference in filter widths: the H and filters are much broader than our F162M and F212N filters, leading to significantly different transmission curves. Besides, these wide-band filters are affected by non-linear effects.
In addition, our extinction curve is in good agreement with those of Fritz et al. (2011), Nishiyama et al. (2009), Schödel et al. (2010), and Hosek et al. (2019), which were derived for different lines of sight in the GC. This suggests that the extinction index in the inner Galaxy does not vary significantly between different lines of sight within the uncertainties. In contrast, our extinction ratios and indices are not consistent with the steeper curves reported by Alonso-García et al. (2017), Nogueras-Lara et al. (2018), and Nogueras-Lara et al. (2020). We conclude that this discrepancy is most likely due to differences in the filter sets used: in those three studies, the filters are considerably broader than those used here.
We also compared our extinction curve with curves derived from regions outside the GC. We found good agreement with the curve of Indebetouw et al. (2005), supporting the idea of a largely spatially invariant infrared extinction curve, even beyond the GC. Additionally, our extinction curve agrees with those of Román-Zúñiga et al. (2007) and Damineli et al. (2016), but only in the SW regime. Moreover, it is consistent with the curve reported by Fahrion & De Marchi (2023) for 30 Doradus, although the uncertainties in that study are too large to allow for definitive conclusions.
@@ Acknowledgements.
LBF and RS acknowledge support from grant PID2022-661136640NB-C21, funded by MCIN/AEI/10.13039/501100011033 and by the European Union. FNL acknowledges support from grant PID2024-162148NA-I00, funded by MCIN/AEI/10.13039/501100011033 and the European Regional Development Fund (ERDF) “A way of making Europe”. LBF, FNL, and RS acknowledge support from the Severo Ochoa grant CEX2021-001131-S, funded by MCIN/AEI/10.13039/501100011033. R.F. acknowledges support from the grants PID2023-146295NB-I00, and from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033 and by “European Union NextGenerationEU/PRTR”. Y.C. was partially supported by a Grant-in-Aid for Scientific Research (KAKENHI number JP24K17103) of the JSPS.References
- Akaike (1974) Akaike, H. 1974, IEEE Transactions on Automatic Control, 19, 716
- Alonso-García et al. (2017) Alonso-García, J., Minniti, D., Catelan, M., et al. 2017, ApJ, 849, L13
- Astropy Collaboration et al. (2022) Astropy Collaboration, Price-Whelan, A. M., Lim, P. L., et al. 2022, ApJ, 935, 167
- Astropy Collaboration et al. (2018) Astropy Collaboration, Price-Whelan, A. M., Sipőcz, B. M., et al. 2018, AJ, 156, 123
- Astropy Collaboration et al. (2013) Astropy Collaboration, Robitaille, T. P., Tollerud, E. J., et al. 2013, A&A, 558, A33
- Bally et al. (2025) Bally, J., Crowe, S., Fedriani, R., et al. 2025, ApJ, 983, 20
- Burriesci (2005) Burriesci, L. G. 2005, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 5904, Cryogenic Optical Systems and Instruments XI, ed. J. B. Heaney & L. G. Burriesci, 21–29
- Bushouse et al. (2023) Bushouse, H., Eisenhamer, J., Dencheva, N., et al. 2023, JWST Calibration Pipeline
- Chiar & Tielens (2006) Chiar, J. E. & Tielens, A. G. G. M. 2006, ApJ, 637, 774
- Chu et al. (2020) Chu, L. E. U., Hodapp, K., & Boogert, A. 2020, ApJ, 904, 86
- Chu & Hodapp (2021) Chu, L. E. U. & Hodapp, K. W. 2021, ApJ, 918, 2
- Crowe et al. (2025) Crowe, S., Fedriani, R., Tan, J. C., et al. 2025, ApJ, 983, 19
- Damineli et al. (2016) Damineli, A., Almeida, L. A., Blum, R. D., et al. 2016, MNRAS, 463, 2653
- de Graauw et al. (1996) de Graauw, T., Haser, L. N., Beintema, D. A., et al. 1996, A&A, 315, L49
- Diolaiti et al. (2000) Diolaiti, E., Bendinelli, O., Bonaccini, D., et al. 2000, A&AS, 147, 335
- Draine (2003) Draine, B. T. 2003, ARA&A, 41, 241
- Eisenhauer et al. (2003) Eisenhauer, F., Abuter, R., Bickert, K., et al. 2003, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4841, Instrument Design and Performance for Optical/Infrared Ground-based Telescopes, ed. M. Iye & A. F. M. Moorwood, 1548–1561
- Fahrion & De Marchi (2023) Fahrion, K. & De Marchi, G. 2023, A&A, 671, L14
- Fritz et al. (2011) Fritz, T. K., Gillessen, S., Dodds-Eden, K., et al. 2011, ApJ, 737, 73
- Gallego-Cano et al. (2018) Gallego-Cano, E., Schödel, R., Dong, H., et al. 2018, A&A, 609, A26
- Genzel et al. (2010) Genzel, R., Eisenhauer, F., & Gillessen, S. 2010, Reviews of Modern Physics, 82, 3121
- Ginsburg et al. (2023) Ginsburg, A., Barnes, A. T., Battersby, C. D., et al. 2023, ApJ, 959, 36
- Girardi (2016) Girardi, L. 2016, ARA&A, 54, 95
- Gordon (2024) Gordon, K. 2024, The Journal of Open Source Software, 9, 7023
- Guarcello et al. (2025) Guarcello, M. G., Almendros-Abad, V., Lovell, J. B., et al. 2025, A&A, 693, A120
- Henshaw et al. (2023) Henshaw, J. D., Barnes, A. T., Battersby, C., et al. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 83
- Hensley & Draine (2023) Hensley, B. S. & Draine, B. T. 2023, ApJ, 948, 55
- Hosek et al. (2018) Hosek, Jr., M. W., Lu, J. R., Anderson, J., et al. 2018, ApJ, 855, 13
- Hosek et al. (2019) Hosek, Jr., M. W., Lu, J. R., Anderson, J., et al. 2019, ApJ, 870, 44
- Indebetouw et al. (2005) Indebetouw, R., Mathis, J. S., Babler, B. L., et al. 2005, ApJ, 619, 931
- Jones & Hyland (1980) Jones, T. J. & Hyland, A. R. 1980, MNRAS, 192, 359
- Kendrew et al. (2013) Kendrew, S., Ginsburg, A., Johnston, K., et al. 2013, ApJ, 775, L50
- Kissler-Patig et al. (2008) Kissler-Patig, M., Pirard, J. F., Casali, M., et al. 2008, A&A, 491, 941
- Lang et al. (2010) Lang, C. C., Goss, W. M., Cyganowski, C., & Clubb, K. I. 2010, ApJS, 191, 275
- Lutz (1999) Lutz, D. 1999, in ESA Special Publication, Vol. 427, The Universe as Seen by ISO, ed. P. Cox & M. Kessler, 623
- Maíz Apellániz et al. (2020) Maíz Apellániz, J., Pantaleoni González, M., Barbá, R. H., García-Lario, P., & Nogueras-Lara, F. 2020, MNRAS, 496, 4951
- Moultaka et al. (2009) Moultaka, J., Eckart, A., & Schödel, R. 2009, ApJ, 703, 1635
- Nishiyama et al. (2006) Nishiyama, S., Nagata, T., Kusakabe, N., et al. 2006, ApJ, 638, 839
- Nishiyama et al. (2008) Nishiyama, S., Nagata, T., Tamura, M., et al. 2008, ApJ, 680, 1174
- Nishiyama et al. (2009) Nishiyama, S., Tamura, M., Hatano, H., et al. 2009, ApJ, 696, 1407
- Nogueras-Lara (2024) Nogueras-Lara, F. 2024, A&A, 681, L21
- Nogueras-Lara et al. (2018) Nogueras-Lara, F., Gallego-Calvente, A. T., Dong, H., et al. 2018, A&A, 610, A83
- Nogueras-Lara et al. (2019a) Nogueras-Lara, F., Schödel, R., Gallego-Calvente, A. T., et al. 2019a, A&A, 631, A20
- Nogueras-Lara et al. (2019b) Nogueras-Lara, F., Schödel, R., Najarro, F., et al. 2019b, A&A, 630, L3
- Nogueras-Lara et al. (2021a) Nogueras-Lara, F., Schödel, R., & Neumayer, N. 2021a, A&A, 653, A33
- Nogueras-Lara et al. (2021b) Nogueras-Lara, F., Schödel, R., & Neumayer, N. 2021b, A&A, 653, A133
- Nogueras-Lara et al. (2020) Nogueras-Lara, F., Schödel, R., Neumayer, N., et al. 2020, A&A, 641, A141
- Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011, Journal of Machine Learning Research, 12, 2825
- Riaz et al. (2009) Riaz, B., Martín, E. L., Bouy, H., & Tata, R. 2009, ApJ, 700, 1541
- Rieke et al. (2005) Rieke, M. J., Kelly, D., & Horner, S. 2005, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 5904, Cryogenic Optical Systems and Instruments XI, ed. J. B. Heaney & L. G. Burriesci, 1–8
- Rodrigo & Solano (2020) Rodrigo, C. & Solano, E. 2020, in XIV.0 Scientific Meeting (virtual) of the Spanish Astronomical Society, 182
- Rodrigo et al. (2012) Rodrigo, C., Solano, E., & Bayo, A. 2012, SVO Filter Profile Service Version 1.0, IVOA Working Draft 15 October 2012
- Román-Zúñiga et al. (2007) Román-Zúñiga, C. G., Lada, C. J., Muench, A., & Alves, J. F. 2007, ApJ, 664, 357
- Schödel (2010) Schödel, R. 2010, A&A, 509, A58
- Schödel et al. (2014) Schödel, R., Feldmeier, A., Neumayer, N., Meyer, L., & Yelda, S. 2014, Classical and Quantum Gravity, 31, 244007
- Schödel et al. (2010) Schödel, R., Najarro, F., Muzic, K., & Eckart, A. 2010, A&A, 511, A18
- Schwarz (1978) Schwarz, G. 1978, Annals of Statistics, 6, 461
- Scoville et al. (2003) Scoville, N. Z., Stolovy, S. R., Rieke, M., Christopher, M., & Yusef-Zadeh, F. 2003, ApJ, 594, 294
- Straižys & Lazauskaitė (2008) Straižys, V. & Lazauskaitė, R. 2008, Baltic Astronomy, 17, 277
- Tokunaga & Vacca (2005) Tokunaga, A. T. & Vacca, W. D. 2005, PASP, 117, 421
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
Appendix A Effective wavelength
We computed the effective wavelengths following the approach described in Appendix B of Nogueras-Lara et al. (2018). That is, we applied Eq. A3 from Tokunaga & Vacca (2005). We used the transmission curves for the JWST-NIRCam filters (see Sect. 2), available on the instrument website555https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-instrumentation/nircam-filters. We computed for RC stars using the most adequate Kurucz model. We adopted the nuclear stellar disc west value of from Nogueras-Lara et al. (2020), as Sgr C is located at the western edge of the GC. To calculate we used an iterative approach. Namely, we computed a first value for using the relation
@@ Appendix A Effective wavelength
The value of , using the uncertainty given by Nogueras-Lara et al. (2020) ().
@@ Appendix B RC selection boxes
Figures 8 and 9 show the RC selection boxes for the reference and consecutive filter methods, respectively, along with the bins defined for each CMD and the two RC features identified with the GMM. Note that the F162M vs. F115W-F162M and F182M vs. F162M-F182M CMDs are identical in both methods.
Appendix C Foreground population in LW filters CMDs
Figure 11 shows the foreground population selection in the F212N vs. F162M-F212N CMD (top panel) along with the F405N vs. F360M-F405N CMD with the foreground stars in blue (bottom panel). We cut the latter CMD at F405N mag, because that is the saturation limit.

### v2

@@ 5.3.2 Other extinction curves
To compare our extinction curve with others derived from regions outside the GC, we selected the extinction curves reported by Indebetouw2005, RomanZuniga2007, Damineli2016, and Fahrion2023, as they represent a variety of environments. Indebetouw2005 derived their extinction curve using the Two Micron All Sky Survey (2MASS; Skrutskie2006) photometry and Spitzer data in the 3–8 m range, targeting two lines of sight: (a quiescent region) and (the giant H II region RCW 49). RomanZuniga2007 obtained their extinction curve for the 1.25–7.76 m range using ground-based observations with ESO New Technology Telescope (NTT) and VLT, as well as Spitzer-IRAC data, focusing on Barnard 59, a dense, star-forming core in the Pipe Nebula. Damineli2016 combined photometric and spectroscopic data from multiple instruments to derive the extinction curve for Westerlund 1 over the 0.4–4.8 m range. Finally, Fahrion2023 derived the extinction curve for 30 Doradus in the Large Magellanic Cloud using JWST-NIRCam observations in the 0.3–4.8 m range.
@@
The resulting comparison is shown in Fig. 7, where our extinction curve is plotted alongside those from the literature and several theoretical dust models. For the left and centre panels, we included the J-band extinction ratio by interpolating our extinction curve. For the left panel, we adopted the isophotal wavelengths of 2MASS and IRAC filters convolved with a K2 III stellar spectrum, following the approach in Indebetouw2005 and RomanZuniga2007. For the right panel, we used the J-band effective wavelength provided by Damineli2016. The dust grain models shown in the figure are included for reference only, and a detailed comparison is beyond the scope of this work.
@@
Table 5 presents the interpolated extinction ratios for the literature extinction curves, alongside those derived in this work, allowing for a more direct comparison. As seen in both the left panel of Fig. 5 and the table, our results are consistent with those reported by Indebetouw2005. This supports the hypothesis that the infrared extinction curve is independent of the line of sight, as their curve was derived from two distinct Galactic directions. Our extinction ratios also match those reported by RomanZuniga2007, although only in the SW regime. Both studies used Spitzer-IRAC data in the LW range; the discrepancies with our results may be due to the broad Spitzer filter profiles and the fact that Barnard 59 is a cold cloud, where the presence of ices can significantly affect extinction in this regime (Riaz2009; Chu2020; Chu2021). Additionally, we computed the ratio using our extinction curve, as described in Sect. 5.3.1, obtaining a value of . This is higher than the values reported by Indebetouw2005 and RomanZuniga2007.
@@
Our extinction curve also shows agreement with that reported by Damineli2016 in the SW regime—except for the ratio—but deviates in the LW regime, as illustrated in the right panel of Fig. 7 and in Table 5. However, a robust comparison is limited by the fact that Damineli2016 do not provide uncertainties for their extinction ratios. Unlike the B59 case, this discrepancy is unlikely to be driven by ice absorption, as the line of sight towards Westerlund 1 shows little evidence of ice (Guarcello2025). Similarly, our extinction curve agrees with that of Fahrion2023, but the associated uncertainties are too large for firm conclusions to be drawn.
@@
While detailed modelling of the dust properties in Sgr C is beyond the scope of this work, we note that the theoretical curve corresponding to from Draine2003 appears to be a closer match to our data than the model. This is consistent with expectations for denser environments with larger grains. Our results differ from the Hensley2023 model in the LW regime. However, we caution that our extinction curve is derived from a localised region, and no strong conclusions about dust grain properties or can be drawn.
@@
In summary, our extinction curve follows the same general trend as the selected literature curves derived outside the GC. It shows excellent agreement with the curve by Indebetouw2005, reinforcing the idea of a broadly line-of-sight–independent infrared extinction curve. It is also consistent with the curve by Fahrion2023, although due to the large uncertainties we cannot perform a definitive assessment.
@@ 6 Conclusions
In this work we aimed to determine the extinction curve towards the Sgr C region using JWST-NIRCam data, and to evaluate its consistency with previous results. We applied two different methods based on the slopes of the RC features to derive the extinction ratios. The results are shown in Tables 2 and 3.
We find that the extinction curve flattens in the LW regime (3.6–4.8 m), in agreement with previous literature. Although we obtained different extinction indices for the SW and LW regimes, we found no significant wavelength dependence within either regime. However, using GALACTICNUCLEUS data overlapping with our JWST field, we derived an H– extinction index consistent with the results of NoguerasLara2020 but not with our own results. We attribute this discrepancy mainly to the difference in filter widths: the H and filters are much broader than our F162M and F212N filters, leading to significantly different transmission curves. Additionally, these wide-band filters are affected by non-linear effects.
Furthermore, our extinction curve is in good agreement with those of Fritz2011, Nishiyama2009, Schoedel2010, and Hosek2019, which were derived for different lines of sight in the GC. This suggests that the extinction index in the inner Galaxy does not vary significantly between different lines of sight within the uncertainties. In contrast, our extinction ratios and indices are not consistent with the steeper curves reported by AlonsoGarcia2017, NoguerasLara2018, and NoguerasLara2020. We conclude that this discrepancy is most likely due to differences in the filter sets used: in those three studies, the filters are considerably broader than those used here.
We also compared our extinction curve with curves derived from regions outside the GC. We find good agreement with the curve of Indebetouw2005, supporting the idea of a largely spatially invariant infrared extinction curve, even beyond the GC. Additionally, our extinction curve agrees with those of RomanZuniga2007 and Damineli2016, but only in the SW regime. Moreover, it is consistent with the curve reported by Fahrion2023 for 30 Doradus, although the uncertainties in that study are too large for definitive conclusions to be drawn.
@@ Acknowledgements.
This work is based on observations made with NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program 4147. The specific observations analysed can be accessed via https://doi.org/10.17909/pwkh-ec13. LBF and RS acknowledge support from grant PID2022-661136640NB-C21, funded by MCIN/AEI/10.13039/501100011033 and by the European Union. FNL acknowledges support from grant PID2024-162148NA-I00, funded by MCIN/AEI/10.13039/501100011033 and the European Regional Development Fund (ERDF) “A way of making Europe”. LBF, FNL, and RS acknowledge support from the Severo Ochoa grant CEX2021-001131-S, funded by MCIN/AEI/10.13039/501100011033. R.F. acknowledges support from the grants PID2023-146295NB-I00, and from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033 and by “European Union NextGenerationEU/PRTR”. Y.C. was partially supported by a Grant-in-Aid for Scientific Research (KAKENHI number JP24K17103) of the JSPS. ZYL is supported in part by NASA 80NSSC20K0533 and NSF AST-2307199.Appendix A Effective wavelength
We computed the effective wavelengths following the approach described in Appendix B of NoguerasLara2018. That is, we applied Eq. A3 from Tokunaga2005. We used the transmission curves for the JWST-NIRCam filters (see Sect. 2), available on the instrument website111111https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-instrumentation/nircam-filters. We computed for RC stars using the most adequate Kurucz model. We adopted the nuclear stellar disc west value of from NoguerasLara2020, as Sgr C is located at the western edge of the GC. To calculate we used an iterative approach. Namely, we computed a first value for using the relation
@@
The value of , using the uncertainty given by NoguerasLara2020.
@@ Appendix B RC selection boxes
Figures 8 and 9 show the RC selection boxes for the reference and consecutive filter methods, respectively, along with the bins defined for each CMD and the two RC features identified with the GMM. Note that the F162M versus F115W-F162M and F182M versus F162M-F182M CMDs are identical in the two methods.
Appendix C Foreground population in LW filter CMDs
Figure 11 shows the foreground population selection in the F212N versus F162M-F212N CMD (top panel) along with the F405N versus F360M-F405N CMD with the foreground stars in blue (bottom panel). We cut the latter CMD at F405N = 12 mag because that is the saturation limit.